$2000 invested at 8% for 30 years what would the return be? What would be the total for 26 years?
Answer by stanbon(75887)

$2000 invested at 8% compounded yearly for 30 years.
what would the return be?
---
A(8) = 2000(1+0.08)^(8) = $3701.86
-----------------------------------------
What would be the total for 26 years?
A(26) = 2000(1.08)^26 = $14,792.71
=========================================
Cheers,
Stan H.
===========